# Measuring Social Biases in Masked Language Models by Proxy of Prediction Quality

### v1

@@ 3.1.1 Benchmark Datasets for Social Bias
The CPS and SS datasets contain biased sentences towards advantaged and disadvantaged groups, where CPS sentence pairs are categorized by bias types: race, age, socioeconomic, disability, religion, physical appearance, gender, sexual orientation and nationality, and SS sentence pairs are categorized by bias types: race, religion, gender and profession. To probe for biases of interest that are encoded in MLMs under consideration, the scope of our experiments include all bias categories and sentences pairs in CPS and intrasentence pairs in SS, since intersentence pairs are not masked for bias evaluation. [^2] Our reported results on the SS dataset include only intrasentence pairs as in experiment code used by Kaneko and Bollegala, 2022. We attempt to estimate the preference of an MLM towards a stereotypical sentence over a less stereotypical one for each bias category in CPS and SS and report corresponding results.
@@
Race
516
Religion
105
Nationality
159
Socioecnomic
172
Gender
262
Sexual orientation
84
Age
87
Disability
60
Physical appearance
63
@@
Race
962
Religion
79
Gender
255
Profession
810
@@
Re-training Dataset
For this work, we focus on recovering encoded social biases from MLMs re-trained on $S^{\text{dis}}$ or $S^{\text{adv}}$ ($\forall s\in S^{\text{dis}}$ or $\forall s\in S^{\text{adv}}$, where $s$ is a sentence) in CPS, to demonstrate the sensitivity of our proposed measures compared to others for characterizing biases introduced by re-training MLMs, but emphasize the substantive use cases of applying our proposed framework beyond a particular corpus type. Biases widely acknowledged in the United States are well represented in CPS [^3] CPS categories are a ”narrowed” version of the US Equal Employment Opportunities Commission’s list of protected categories. compared to SS, and there is greater diversity of sentence structures in CPS Nangia et al. (2020).
The validation rate of CPS is 18% higher than SS, and paired with concerns regarding standard masked language modeling metric viability on SS, we proceed to use sentence sets in CPS as opposed to SS to re-train MLMs (and validate methods for estimating biases introduced by that re-training), given that CPS is a more reliable benchmark for pre-trained MLM bias measurement Nangia et al. (2020).
$\forall s\in S^{\text{dis}}$ for CPS
Uncased Cased
Unique Tokens
4631
4800
Lexical Diversity
0.235
0.244
$\forall s\in S^{\text{adv}}$ for CPS
Uncased Cased
Unique Tokens
4607
4768
Lexical Diversity
0.233
0.241
Table 2: Data profile for (cased and uncased) sentence sets used to re-train MLMs, where lines correspond to sentences in CPS as detailed in 3.1.1. Includes 1508 lines and total tokens greater than $1.9*10^{4}$ for both cased and uncased data.
3.2 MLM Preference by Prediction Quality
When considering a sentence $s$ containing tokens $\{t_{1},t_{2},...,t_{l_{s}}\}$, where $l_{s}$ is the number of tokens in $s$, (modified) token(s) of $s$ can characterize its bias towards either disadvantaged or advantaged groups. For example, in the sentence pair $s_{1}\in S^{dis}$ and $s_{2}\in S^{adv}$ from the CPS dataset, where $s_{1}$ is "The poor are really ignorant about how to handle the money they do have" and $s_{2}$ is "The rich are really ignorant about how to handle the money they do have", the bias is characterized by modified tokens "poor" and "rich" respectively. All other tokens between $s_{1}$ and $s_{2}$ in this example are unmodified.
For a given sentence $s$ with tokens $t\in s$ we denote all tokens besides $t_{x}$ as $s_{\setminus t_{x}}$ (where $1\leq x\leq l_{s}$). For a given sentence $s$, we denote modified tokens as $M$ and unmodified tokens as $U$ ($s=U\cup M)$. For a given MLM with parameters $\theta$, we denote a masked token as $t_{m}$ and a predicted token as $t_{p}$.

### v2

@@ 3.1 Benchmark Datasets for Social Bias
CPS and SS contain biased sentences towards advantaged and disadvantaged groups (see Appendix A for specific bias categories and sentence counts). To probe for biases of interest that are encoded in considered MLMs, the scope of our experiments include all bias categories and sentences pairs in CPS and intrasentence pairs in SS, since intersentence pairs are not masked for bias evaluation (Kaneko and Bollegala, 2022). We estimate the preference of an MLM towards a stereotypical sentence over a less stereotypical one for each bias category in CPS and SS and report corresponding results.
3.2 Retraining Dataset
CPS provides a more diverse alternative to biases expressed by sentence pairs in SS. Biases widely acknowledged in the United States are well represented in CPS [^3] CPS categories are a ”narrowed” version of the US Equal Employment Opportunities Commission’s list of protected categories Nangia et al. (2020). compared to SS, and there is greater diversity of sentence structures in CPS Nangia et al. (2020). CPS has been found to be a more reliable benchmark for pretrained MLM bias measurement than SS, and the validation rate of CPS is 18% higher than SS. Nangia et al. (2020). Based on these findings, paired with (1) the computational expense and time-consumption involved with retraining MLMs under the MLMO and (2) concerns regarding standard masked language modeling metric viability on SS, we proceed to use sentence sets in CPS to re-train MLMs and validate methods for estimating the biases that are introduced Nangia et al. (2020). The data profile for (cased and uncased) sentence sets used to re-train MLMs is shown in Appendix A.2. We re-train MLMs on $\forall s\in S^{\text{dis}}$ or $\forall s\in S^{\text{adv}}$, where $s$ is a sentence biased towards either advantaged ($S^{\text{adv}}$) or disadvantaged groups ($S^{\text{dis}}$), and compare the accuracy of our proposed measures with others.
3.3 Transformer-based Language Models
We denote re- and pretrained transformers as $T_{R}$ and $T_{P}$ respectively. The subscript ${}_{\text{unc}}$ denotes an uncased model. We report and compare results from the following transformer-based language models available through the HuggingFace library Wolf et al. (2020): bert-base-uncased ($\text{BERT}_{\text{unc}}$; Devlin et al., 2019), roberta-base (RoBERTa; Liu et al., 2019b), distilbert-base-uncased ($\text{distilBERT}_{\text{unc}}$; Sanh et al., 2019), and distilroberta-base (distilRoBERTa; Liu et al., 2019a).
3.4 Recovering Social Biases in Pretrained MLMs
Model ime Input with masked token $w_{11}$ = money $w_{11}$ Rank
$\text{distilRoBERTa}_{P}$ The poor are really ignorant about how to handle the problems they do have 5
$\text{distilRoBERTa}_{R}$ The poor are really ignorant about how to handle the money they do have 1
$\text{distilRoBERTa}_{P}$ The rich are really ignorant about how to handle the wealth they do have 4
$\text{distilRoBERTa}_{R}$ The rich are really ignorant about how to handle the problems they do have 2
Table 1: MLM predictions and ground truth masked token ranks for input sentences under the ime. The processed input sentences are "The [poor, rich] are really ignorant about how to handle the money they do have", and are contained in $S^{\text{dis}}$ or $S^{\text{adv}}$ from CPS with characteristic words poor and rich respectively. Masked tokens are in bold.
We probe for MLM preferences using the ime, which masks one token at a time until all tokens have been masked, or until we have $n$ logits or predictions for a sentence with $n$ tokens. Appendix I shows an example of the ime for one model and text example.
Table 1 shows masked token predictions (those with first rank and highest probability) produced by MLMs for example input contexts. The ground truth masked token rank, or $\rho(t_{m}|c)$, represents the quality of the models’ predictions relative to a sentence biased against an advantaged or disadvantaged group. Special start and end character tokens for MLMs are not included in the span of tokens considered in our experiments to eliminate noise. [^4] Special start and end character tokens for MLMs are not considered by measures using attention weights and probabilities computed from the (log)softmax of model logits for the masked token index.
For a given MLM with parameters $\theta$, we propose attention-weighted measures $\textsc{crra}(w)$ and $\Delta\text{PA}(w)$ defined in Equations 4 and 5 respectively, where $a_{w}$ is the average of all multi-head attentions associated with the ground truth token $w$, and $P(t_{m}|s_{\setminus t_{m}})$ and $\rho(t_{m}|s_{\setminus t_{m}})$ are the probability score and rank of the masked token respectively. We extend these definitions for a sentence $s$ as shown in Equations 6 and 7.
$\textsc{crra}(s):=\frac{1}{l_{s}}\sum_{w\in s}\textsc{crra}(w;\theta)$ (6)
We compute measure $f$, $\forall f\in\{\textsc{crr}_{T}(s),\\ \Delta\textsc{p}_{T}(s),\textsc{crra}_{T}(s),\Delta\textsc{pa}_{T}(s)\}$, where $T$ is an MLM transformer and $s$ is a sentence, $\forall s\in S^{\text{dis}}$ and $\forall s\in S^{\text{adv}}$. [^5] We apply the Shapiro-Wilk test Shapiro and Wilk (1965) for normality to each measure and did not find evidence that the measures were not drawn from a normal distribution. The same was found for the difference of each of these measures between sentence sets relative to the same transformer $T$. Appendix N shows these measures (likelihood scores) for an example sentence $s$ in SS and CPS. [^6] Greater MLM preference based on prediction quality is reflected by crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ values closer to 0 (if a sentence with bias against advantaged groups has a greater value relative to its paired counterpart, the MLM is deemed to prefer bias against disadvantaged groups). The opposite is true for measures csps, sss, aul and aula.
We define sets of measures $M_{1}$ and $M_{2}$ ($M_{1}\cap M_{2}=\emptyset$), where $M_{1}=\{\Delta\textsc{p},\textsc{crr},\textsc{crra},\Delta\textsc{pa}\}$ and $M_{2}=\{\textsc{aul},\textsc{aula},\textsc{csps},\textsc{sss}\}$.
We apply Equation 8 to estimate the preference of a transformer $T$ for $s\in S^{\text{dis}}$ relative to $s\in S^{\text{adv}}$ for paired sentence $s$ and measure $f$, $\forall f\in\{\textsc{crr},\\ \textsc{crra},\Delta\textsc{p},\Delta\textsc{pa},\text{{csps}},\textsc{sss},\text{{aul}},\text{{aula}}\}$. We define a bias score for a pretrained MLM as bspt (Equation 9).
$\mathds{1}$ is a per-model indicator function which returns $1$ if transformer $T$ has a larger preference for $S^{\text{dis}}$ relative to $S^{\text{adv}}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bspt represents the proportion of sentences with higher relative bias against disadvantaged groups for a given measure, where values above 50 indicate greater relative bias against disadvantaged groups for an MLM.
3.4.1 Pretrained MLM Bias Scores
We use bspt to compare preferences for a given MLM and report results for all considered measures and MLMs in Appendix K, including corresponding tables and a detailed analysis of results for CPS and SS, focusing on race bias in particular. We report overall bias scores in Table 2.
[…]
CPS Dataset
csps 59.35 60.48 59.35 56.83
aul 58.75 48.34 53.32 51.59
aula 58.09 48.21 51.86 52.65
crr 58.89 61.07 57.76 56.23
crra 60.68 58.89 61.94 60.08
$\Delta\textsc{p}$ 59.88 60.08 59.75 57.49
$\Delta\textsc{pa}$ 60.15 60.81 59.81 58.02
SS Dataset
sss 61.06 59.16 61.4 60.59
aul 59.45 48.91 60.21 51.71
aula 58.83 50.28 59.59 51.66
crr 57.83 53.85 54.37 53.42
crra 62.06 58.59 60.54 61.11
$\Delta\textsc{p}$ 62.2 58.64 61.4 59.31
$\Delta\textsc{pa}$ 62.35 58.21 61.35 59.31
Table 2: Overall bias scores for pretrained MLMs using bspt$(f)$ with considered measures $\forall f\in\{\textsc{crr}_{T}(s),\\ \Delta\textsc{p}_{T}(s),\textsc{crra}_{T}(s),\Delta\textsc{pa}_{T}(s)\}$ on CPS and SS, where D- denotes retrained. Color-coded from darkest to lightest, with lower values represented by the darkest shade and higher values by the lightest shade.
Overall, all evaluation methods show concerning social biases against disadvantaged groups embedded in MLMs as observed in prior research (Kaneko and Bollegala, 2022, Nangia et al., 2020). Interestingly, $\text{BERT}_{\text{unc}}$ has the lowest overall sss, aul, aula, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ (second lowest crr) on SS, but conflicting results on CPS, where it has the highest csps, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ but the lowest aul, aula, and crra. RoBERTa and distilRoBERTa have higher overall bias than $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ according to (1) all but one measure on SS and (2) aul and crra on CPS.
[…]
We find that only aul and aula estimate overall bias scores below 50, where the corresponding MLM is $\text{BERT}_{\text{unc}}$ in each case.
3.4.2 Alignment with Human Annotated Biases
We compare the alignment (agreement) between measures and bias ratings in CPS. Sentence pairs in CPS received five annotations in addition to the implicit annotation from the writer. We map these sentences to a binary classification task, where a sentence is considered biased if it satisfies criteria from Nangia et al., 2020, where (1) at least three out of six annotators (including the implicit annotation) agree a given pair is socially biased and (2) the majority of annotators who agree a given pair is socially biased agree on the type of social bias being expressed. [^9] This experiment setting gives 58 unbiased pairs and 1,450 biased pairs for binary classification.
We compute evaluation measures derived from MLMs to predict whether a pair is biased or unbiased at varying thresholds. All measures are computed for each sentence in a pair. Thresholds for bias scores computed on sentences with bias against advantaged and disadvantaged groups respectively maximize area under the ROC Curve for that measure. We find that one or more of our measures outperform aul and aula in their agreement with human annotators on CPS for $\text{RoBERTa}_{P}$ and $\text{BERT}_{P,\text{unc}}$ based on higher AUROC values if considered MLMs exhibit bias towards disadvantaged groups (ROC curves shown in Appendix M).
3.5 Recovering Social Biases Introduced by Retraining MLMs
We re-train each transformer under consideration using the PyTorch Python library with P100 and T4 GPUs on cased (RoBERTa and distilRoBERTa) and uncased ($\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$) versions of CPS sentences. [^10] The results are from retraining with 0.15 mlm probability (as in Devlin et al., 2019), using 80 percent of data for training and 20 percent for validation. MLMs were retrained for 30 epochs, reaching minimum validation loss at epoch 30.
Figure 1: Difference between ECDFs for crr distribution for sentences in $S^{\text{dis}}$ and $S^{\text{adv}}$ for re- and pretrained RoBERTa and the race bias category in CPS. The line at $y=0$ separates what is biased against disadvantaged groups on the positive y-axis from what is biased against advantaged groups on the negative y-axis.
Figure 1 shows the difference in ECDFs for the $\textsc{crr}_{m}(s)$ measure on the race bias category in CPS, where $m$ is retrained MLM $\text{RoBERTa}_{R}$. When compared with pretrained RoBERTa (RoBERTa-P), we observe an upwards shift in the difference of ECDFs for the crr difference between sentence sets after retraining RoBERTa on $S^{\text{dis}}$ (RoBERTa-R-DIS), and a downwards shift after retraining on $S^{\text{adv}}$ (RoBERTa-R-ADV). This is expected since retraining MLMs on $S^{\text{dis}}$ or $S^{\text{adv}}$ should shift the MLM preference towards the corresponding bias type. In this sense, the figure illustrates how the difference in ECDF distributions for an MLM can visually represent a contextual shift in relative bias using proposed measure crr.
We define a bias score for a retrained MLM (relative to its pretrained base) as bsrt (Equation 10).
We define a proxy for the relative preference between two MLMs with the model-comparative indicator function $\mathds{1}$, which returns $1$ if transformer $T_{1}$ has a larger preference for $S^{\text{dis}}$ relative to $S^{\text{adv}}$ than transformer $T_{2}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bsrt can be applied to compare pre- and retrained MLMs within the same model class and recover biases introduced by MLM retraining. Values above 50 indicate greater bias against disadvantaged groups for transformer $T_{1}$ relative to $T_{2}$.
3.5.1 Retrained MLM Bias Scores
We compute bsrt for retrained MLMs and all measures and bias categories on CPS and report results in Appendix P. [^11] We assess whether the relative differences in proportions of bias between re- and pretrained transformers are significant according to McNemar’s test McNemar (1947). In general, each measure produces a bias score in accordance with the particular retraining dataset used ($S^{\text{dis}}$ or $S^{\text{adv}}$) for almost all significant results across MLMs, demonstrating that proposed function bsrt can be applied to estimate the bias against disadvantaged groups for transformer $T_{1}$ relative to its pretrained base $T_{2}$.
We find that crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are typically more accurate than csps, aul, and aula for measuring social biases introduced by retraining MLMs, and outperform others with regard to sensitivity for relative changes in MLM bias due to retraining, indicated by larger and smaller scores and more frequently significant relative difference in proportions of bias between re- and pretrained transformers. In every case and across considered MLMs, one or more of our measures reports the highest retraining bias scores for MLMs retrained on $S_{\text{dis}}$ and the lowest for MLMs retrained on $S_{\text{adv}}$.
For $\text{RoBERTa}_{R}$ retrained on $S_{\text{adv}}$ from CPS, each measure gives scores below 50 as expected. However, aul and aula give insignificant results for physical appearance bias.121212crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give significant results below 50 for each bias category with lower scores than other measures in general. For $\text{BERT}_{R}$ retrained on $S_{\text{adv}}$, aul and aula overestimate physical appearance, gender, disability, and socioeconomic biases and give insignificant results for each, while all proposed measures give significant results below 50 for each category as expected. Similarly, for $\text{distilRoBERTa}_{R}$, aula overestimates physical appearance bias and gives insignificant results for physical appearance, gender, and age, aul gives insignificant results for physical appearance, and csps and aul give insignificant results for age.131313aul and aula also give insignificant results for physical appearance, disability, and sexual orientation for $\text{distilBERT}_{R}$ retrained on $S_{\text{adv}}$. Overall, we find that aul, aula, and csps overestimate physical appearance bias introduced by retraining MLMs compared to proposed measures and based on bsrt.
For all MLMs retrained on $S_{\text{dis}}$, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results above 50 for each bias category as expected, with higher scores than other measures in almost every case. [^14] This also applies to crra except for disability bias using $\text{distilBERT}_{R}$. In addition, proposed measures are significant for every bias type using $\text{BERT}_{R}$ and $\text{RoBERTa}_{R}$. This is also true for $\text{distilBERT}_{R}$ (besides crra for sexual orientation) and $\text{distilRoBERTa}_{R}$ (besides crra for physical appearance and crr for disability). In contrast, aul, aula, and csps have 5, 6, and 11 insignificant results respectively across all considered MLMs retrained on $S_{\text{dis}}$, and aula and csps each give 3 and 4 bias scores below 50. Notably, aul and crra each give 1 bias score below 50. These are concerning underestimations of biases introduced by retraining MLMs only on sentences with biases against disadvantaged groups from CPS.151515csps, and aula for $\text{RoBERTa}_{R}$ (Appendix P) give scores less than 50 for age bias. aula also gives a score less than 50 for sexual orientation bias. csps, aul, and aula give insignificant results for sexual orientation and age, along with religion and physical appearance for csps and disability for aula.
3.5.2 Alignment with Retraining Biases
We frame a binary classification task where bsrt above 50 indicates increased preference (after retraining) for sentences with bias against disadvantaged groups in CPS (1), and vice versa for scores below 50 (0). [^16] There are 72 predictions per measure across MLMs and bias types. Half of binary truths are 1 and 0 respectively since MLMs retrained on $S_{\text{dis}}$ should score above 50 for all bias types (vice versa for MLMs retrained on $S_{\text{adv}}$) and $S_{\text{dis}}$ has the same length (number of sentences) as $S_{\text{adv}}$.
Re-train Dataset: $\forall s\in s^{dis}$ for CPS
Model csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
$\text{BERT}_{R,\text{unc}}$ 0.028 0.028 0.028 0.000 0.000 0.000 0.000
$\text{RoBERTa}_{R}$ 0.028 0.000 0.083 0.000 0.000 0.000 0.000
$\text{distilBERT}_{R,\text{unc}}$ 0.000 0.000 0.000 0.000 0.028 0.000 0.000
$\text{distilRoBERTa}_{R}$ 0.056 0.000 0.000 0.000 0.000 0.000 0.000
Re-train Dataset: $\forall s\in s^{adv}$ for CPS
Model csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
$\text{BERT}_{R,\text{unc}}$ 0.000 0.083 0.083 0.000 0.000 0.000 0.000
$\text{RoBERTa}_{R}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\text{distilBERT}_{R,\text{unc}}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\text{distilRoBERTa}_{R}$ 0.000 0.000 0.028 0.000 0.000 0.000 0.000
Table 3: Error rates for MLMs using considered measures for the binary classification task described in Section 3.5.2. Bold values indicate the lowest error rate for an MLM across all measures.
We report error rates for MLMs in Table 3, and find that one or more of proposed measures produce the lowest error rate for all considered MLMs. crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are 100% accurate and crra is about 99% accurate, while aul, aula, and csps are about 93%, 88%, and 94% accurate respectively. Based on this evaluation setting, proposed measures crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are more accurate than csps, aul, and aula for estimating social biases introduced by retraining MLMs.
4 Conclusion
We represent MLM bias through a model’s relative preference for ground truth tokens between two paired sentences with contrasting social bias under the ime, measuring it using the (attention-weighted) quality of predictions. We evaluate social biases for four state-of-the-art transformers using benchmark datasets CPS and SS and approximate the distributions of proposed measures. We use bspt to compute bias scores for pretrained MLMs using considered measures, and find that all encode concerning social biases. We find that gender has lower encoded biases on CPS compared to SS across MLMs, and that other measures can underestimate bias against disadvantaged groups in the religion category on SS and disability category on CPS. We propose bsrt to estimate the social biases against disadvantaged groups for a retrained MLM relative to its pretrained base, and assess bias evaluation measures for alignment with biases introduced by MLM retraining under the MLMO. We find that proposed measures (such as crra and $\Delta\textsc{pa}$) produce more accurate estimations of biases introduced by MLM retraining than previously proposed ones, which underestimate biases after retraining on sentences biased towards disadvantaged groups. We hope our methods prove useful to the research community for measuring social biases introduced by retraining MLMs.
5 Limitations
We anticipate that the limitations addressed in this section will be useful for future research evaluating social biases in MLMs.
As described in section 3, we leverage sentence pairs from 2 benchmark datasets, CPS and SS, to evaluate the social biases of pre- and retrained MLMs. Both datasets are limited to the English language and specific social bias types represented by binary sentence sets. Future research extending this work could consider and compare alternative benchmark datasets with different languages, social bias types and sentence set structures. In addition, we acknowledge the dependency on human annotated biases in benchmark datasets when assessing discussed measures.
In this work, and as mentioned in section 2.2, we focus on an MLM’s key pretraining objective, masked language modeling, to measure social biases of the MLM. Different pretraining objectives such as next sentence prediction are beyond the scope of this paper. Furthermore, we measure relative changes in biases w.r.t. the intrinsic biases of a base MLM after retraining under the MLMO, and report and compare results from the four transformers mentioned in section 3.3, each of which was retrained for 30 epochs and reached a minimum validation loss at epoch 30. A logical extension of this work would be considering MLMs with different architectures or training data. We propose a model-comparative function bsrt to measure the relative change in MLM biases after retraining. Future research could leverage this function to assess the sensitivity of discussed measures to a range of MLM retraining conditions.
We also encourage research assessing the agreement between relative changes in MLM biases introduced by retraining and biases embedded in the retraining corpus.
6 Ethical Considerations
The methods and measures employed and proposed as part of this work are intended to be used for measuring social biases in pre- and retrained MLMs. We do not condone the use of this research to further target disadvantaged groups in any capacity. Instead, we encourage the use of proposed measures in conjunction with model debiasing efforts to lessen encoded social biases against disadvantaged groups in MLMs used in production settings. No ethical issues have been reported concerning the datasets or measures used in this paper to the best of our knowledge.
Appendix A Datasets
A.1 Benchmark Datasets for Social Biases
@@
Race 516
Religion 105
Nationality 159
Socioecnomic 172
Gender 262
Sexual orientation 84
Age 87
Disability 60
Physical appearance 63
@@
Race 962
Religion 79
Gender 255
Profession 810
@@
A.2 Retraining Datasets
$\forall s\in S^{\text{dis}}$ for CPS unc Cased
Unique Tokens 4631 4800
Lexical Diversity 0.235 0.244
$\forall s\in S^{\text{adv}}$ for CPS unc Cased
Unique Tokens 4607 4768
Lexical Diversity 0.233 0.241
Table 5: Data profile for (cased and uncased) sentence sets used to re-train MLMs, where lines correspond to sentences in CPS as detailed in 3.1. Includes 1508 lines and total tokens greater than $1.9*10^{4}$ for both cased and uncased data.
Appendix B Equation for $\textsc{csps}(s)$